# Eye-tracking-Driven Shared Control for Robotic Arms: Wizard of Oz Studies to Assess Design Choices

### v1

@@ Abstract.
Advances in eye-tracking control for assistive robotic arms provide intuitive interaction opportunities for people with physical disabilities. Shared control has gained interest in recent years by improving user satisfaction through partial automation of robot control. We present an eye-tracking-guided shared control design based on insights from state-of-the-art literature. A Wizard of Oz setup was used in which automation was simulated by an experimenter to evaluate the concept without requiring full implementation. This approach allowed for rapid exploration of user needs and expectations to inform future iterations. Two studies were conducted to assess user experience, identify design challenges, and find improvements to ensure usability and accessibility. The first study involved people with disabilities by providing a survey, and the second study used the Wizard of Oz design in person to gain technical insights, leading to a comprehensive picture of findings.
@@
ccs: Human-centered computing Accessibility systems and toolsccs: Human-centered computing Pointing devicesccs: Human-centered computing Empirical studies in accessibilityccs: Software and its engineering Development frameworks and environments
@@ 1. Introduction
People with physical disabilities such as locked-in syndrome, cerebral palsy, late-stage amyotrophic lateral sclerosis, or multiple sclerosis benefit from non-haptic controls such as Brain-Computer Interfaces and eye-tracking controllers. Implementation as control inputs for assistive robotic arms can enhance user independence, self-determination, and privacy in daily life by reducing the dependency on family members and caregivers (Fischer-Janzen et al., 2024; Awuah et al., 2024; Baumeister et al., 2021).
Regardless of the controlled technology, challenges have been identified with such input devices in the areas of safety, availability of participants for clinical trials, privacy, interdisciplinarity, and cost (Maiseli et al., 2023; Suryadarma et al., 2024; Fischer-Janzen et al., 2024; Nanavati et al., 2024b). This has led to the development of a variety of robot control methods that differ in terms of input modalities and control strategies (Mohebbi, 2020; Fischer-Janzen et al., 2024). Depending on the input modality, the achievable varying level of robot autonomy yields different advantages and disadvantages. For instance, direct control of the robot leads to the need for mode switches to cover the 3-dimensional space the robot can move, or to control all the joints of the robot, enhancing the needed concentration on the tasks (Fischer-Janzen et al., 2024; Kronhardt et al., 2023; Herlant et al., 2016). The implementation of a shared control approach, in which tasks are partially automated, has been demonstrated to effectively address this issue. Such approaches have been shown to positively impact interaction perception and reduce the time required to complete tasks (Bhattacharjee et al., 2020; Kronhardt et al., 2023). A potential challenge associated with a high level of automation is user discomfort, which can arise if users feel that they are losing control over the system. In this context, community feedback is essential to improve design and evaluate usability and accessibility (Stalljann et al., 2020; Baumeister et al., 2021; Nanavati et al., 2024a). It provides firsthand insights into user expectations and perceptions of their daily lives, thus enhancing acceptance rates (Nanavati et al., 2024a; Canal et al., 2021).
This work incorporates feedback from the community on an initial shared control design to identify challenges and improve the design. A Wizard of Oz design of the eye-tracking-driven control was developed, with which two studies were conducted. The first study was an online survey distributed to people with disabilities, their family members and friends, and caregivers. The objective was to assess the usability in daily life and to identify the tasks that participants would prefer to perform. The stakeholder group was defined as individuals with physical disabilities who are unable to use their extremities to control the robot. In the second study, non-disabled participants with varying levels of robotics experience interacted with the system to obtain technical parameters that would provide constraints for a future control algorithm. In this context technical parameters describe for example the maximum and minimum dwell time and the level of accuracy needed for a robust control. The main contributions of this work are as follows:
@@
    Presentation of a shared control design for everyday tasks.
@@
    An evaluation of functionality and usability through opinions of potential users, their families, caregivers and non-disabled participants.
@@
    Improvements in the initial shared control design gained from the studies.
@@ 2. Background
2.1. Shared control for robotic arms
@@
”Shared control in assistive robotics blends human autonomy with computer assistance, thus simplifying complex tasks for individuals with physical impairments” (Goldau et al., 2024). Advances in sensing, inference, modeling, and learning methods have lead to novel developments in the research area (Selvaggio et al., 2021). In this context, researchers are investigating impact factors on shared control. Literature has been reviewed to find such factors influencing human-robot-interaction. three categories were introduced to facilitate the overview over the found outcomes: Field of use, technical challenges, level of autonomy, and user individuality and preferences.
The field of use establishes the boundaries within which shared controls are to be utilized. Robots can be used for daily tasks in unstructured environments (Stanger et al., 1994), household chores (Herr et al., 2016), and personal tasks such as personal hygiene or medical procedures (Wang et al., 2020). This determines the positioning of the robot, whether it must be attached to an electric wheelchair, on the bedside, or on a table (Stanger et al., 1994), and leads to the initial setup configuration and limitation in robot choice.
The technical challenges include reliability and functionality of the system: Safety, robustness and success rate in e.g. task completion is a standard that has to be ensured (Stanger et al., 1994; Wang et al., 2020) and the requirements are defined in ISO 13482 (für Normung e.V., DIN). The system must be robust to changing or unorganized environments, especially when the robot is mounted on an electric wheelchair (Selvaggio et al., 2021; Stanger et al., 1994). The system should be easy to learn and use, which also depends on the complexity of the task (Stanger et al., 1994; Shafti et al., 2019; Goldau et al., 2024). Visualization with an interface can help (Shafti et al., 2019; Nanavati et al., 2024b). If interfaces in the form of displays are not available other feedback methods can ensure the ease of use by communicating information to the user and reducing uncertainty in using the robot (Shafti et al., 2019; Nanavati et al., 2024b). Autonomous robotic task learning, for example imitation learning, ensures the availability of a large range of tasks (Stanger et al., 1994; Selvaggio et al., 2021). The availability of general tasks adapted to the field of usage eases the distress in the lack of capabilities (Brose et al., 2010). The appearance and aesthetics can influence the results of user’s system perception. In the past, connections were found that due to insufficient aesthetics, users expected reduced social communication and isolation (Stanger et al., 1994; Brose et al., 2010).
Ongoing debates revolve around the level of autonomy and can vary from teleoperated systems to fully autonomous robots (Beer et al., 2014). Recent studies have shown that more autonomy is not always better for the user (Nanavati et al., 2024b; Bhattacharjee et al., 2020). Autonomous behavior of the robot ”allows reducing the human operator’s workload when performing a task that can be repetitive and/or requires skills, effort or precision levels that exceed those of a human” (Selvaggio et al., 2021). If the robot produces more errors in task execution with a high level of autonomy, users tend to reject the system (Bhattacharjee et al., 2020). It can further lead to lost trust or a sense of lost control with unpredictable behavior of the system (Nanavati et al., 2024b). Validation functions communicating and verifying the next steps with the user can ease this issue (Bhattacharjee et al., 2020). On the other hand, the user’s attitude towards a task influences the wish for a higher level of autonomy or a complete delegation of the task to the robot, e.g. if the user chooses not to perform the task (Herr et al., 2016). In eye-tracking controls these levels of autonomy can be distinguished in teleoperated system, directionally controlled systems and systems with automated trajectory or grasp planning. This variation strongly influences the design process (Fischer-Janzen et al., 2024). Directional gaze control is easy to learn, provides a sense of control, and reduces cognitive load (Dragomir et al., 2021; Dziemian et al., 2016). On the other hand, it is more time consuming to control each of the robot movements separately. Shared controls are mostly controlled by the user selecting the object via fixation, indicating the wish to interact with (Catalán et al., 2017; Zeng et al., 2017; Tostado et al., 2016).
The user’s individuality and preferences vary with personal metrics such as the user’s age, level of disability, living arrangements, and familiarity with technology and affect the system’s performance perception (Stanger et al., 1994). The adaptation of the system to user preferences is mentioned by various authors (Bhattacharjee et al., 2020; Selvaggio et al., 2021). The ability to adapt the system gave rise to Shared Autonomy approaches, ”where the robot is capable of seamless adaptation of its autonomy level based on its own understanding of the human actions/intentions and of the surrounding environment” (Selvaggio et al., 2021). Incorporating the preferences of the user influences the perception and user satisfaction of the system (Cakmak et al., 2011; Canal et al., 2021). An example is the participant’s preference for handing an object with a particular orientation (Cakmak et al., 2011). A common factor, as known from the use of PCs, is the adaptation of the system to the user’s preferences by offering a settings menu. Parameters such as the velocity of the robot should be adaptable (Bhattacharjee et al., 2020). Another factor is that offering a variation of input modalities (e.g. head or eye-tracking control instead of haptic control) makes it possible to tailor the system to the user’s needs (Bhattacharjee et al., 2020; Nanavati et al., 2024b). Users want their interactions with the robot ”to be customized to their impairment, their preferences, and their context” (Nanavati et al., 2024b).
Other impact factors are a low maintenance system to ensure a reduced workload for others, such as family or caregivers (Bhattacharjee et al., 2020). Related to market introduction are factors such as the initial cost and the maintenance cost of the system (Stanger et al., 1994; Brose et al., 2010). These may have less impact on the outcomes of the user survey, but are an important factor in the design of the system.
Due to the closeness between human and robot these technical and ethical challenges coexist (Canal et al., 2021). In designing a well-balanced shared control, a trade-off must be found for all factors relating to technical and ethical challenges. In the following, we will present a solution that considers and balances the different influences.
2.2. Task preferences in assistive robot manipulation
Among the many daily tasks, research has identified the areas where people with physical disabilities prefer assistance. This can have significant impact on the design of the system, through the adaptation of grippers or, in the case of shared control, task automation.
The four most frequently mentioned tasks that an assistive robot should perform, were food related tasks (serving a meal, eating) (Bien et al., 2003; Stanger et al., 1994; Brose et al., 2010; Chang and Park, 2003; Canal et al., 2021; Bhattacharjee et al., 2020; Nanavati et al., 2024a), drinking related tasks (Bien et al., 2003; Stanger et al., 1994; Brose et al., 2010; Chang and Park, 2003), picking up objects (Bien et al., 2003; Stanger et al., 1994; Brose et al., 2010; Chang and Park, 2003; Nanavati et al., 2024b), and personal hygiene (Stanger et al., 1994; Brose et al., 2010; Nanavati et al., 2024b). Picking up an object was further related to the location or a manipulation that was desired with the object such as picking it up from the floor or from a shelf, bringing it near, or moving it. Personal hygiene included various tasks such as shaving (Bien et al., 2003; Chang and Park, 2003), washing the face and hands, brushing teeth, combing hair, shaving, applying makeup (Brose et al., 2010), and dressing (Canal et al., 2021; Nanavati et al., 2024b).
The importance of being able to manipulate a wide variety of objects benefits the ability to get around, for example with an electric wheelchair by being able to move blocking objects out of the way. This is also a task desired by potential users (Nanavati et al., 2024b).
Miscellaneous tasks were mentioned with taking medicine (Nanavati et al., 2024b), working (Nanavati et al., 2024b), leisure activities (Stanger et al., 1994; Nanavati et al., 2024a), turning on and off switches (Bien et al., 2003; Chang and Park, 2003), opening and closing doors (Bien et al., 2003; Brose et al., 2010; Chang and Park, 2003), scratching oneself (Bien et al., 2003; Chang and Park, 2003) and making tea. As sources from 2003 show, changing CDs and removing paper from fax machines are examples of how task preferences change over time (Bien et al., 2003; Chang and Park, 2003). As the accessibility of smart home applications continues to evolve, it is expected that more tasks will become less important for robot assistance, such as flipping light switches or opening doors in the home.
Based on these results, a first selection of tasks was selected for the Wizard of Oz design to present in the studies. In Section 4.1 the literature findings were used to group the responses from the online survey and to compare results related to the change in task prioritization over time in Section 5.1.
2.3. Gaze as control input
@@
The first eye-tracking controllers for robotic arms were developed as early as 2001 (Kim et al., 2001). Since then, several systems have been developed, of which only a few have become commercially available (Fischer-Janzen et al., 2024; Suryadarma et al., 2024; Brose et al., 2010). They differ in the control approach used to guide the robot. A popular design approach is directional control, where the robot moves in the direction of gaze (Alsharif, 2017; Sunny et al., 2021). Graphical User Interfaces (GUIs) are available in which the robot is moved with gaze-selected buttons, taking advantage of more accurate stationary eye-tracking systems (Dragomir et al., 2021; Stalljann et al., 2020).
Going a step further, human intention in eye-tracking control can be predicted from task-related eye movements (anticipatory gaze (Aronson and Admoni, 2020)) based on the natural gaze behavior in hand-eye coordination (González-Díaz et al., 2024). Task anticipation has been explored to estimate task selection from the gaze pattern (Huang and Mutlu, 2016; Wang, 2021). It can be predicted before the action begins and during its execution (González-Díaz et al., 2024; Selvaggio et al., 2021). A crucial step in using gaze as control input is to decompose the behavior into subtasks to obtain specific gaze patterns that can be used to train AI models (Hayhoe and Ballard, 2014). Another application is to use gaze as a complement to generate more dense and accurate information, as is done in Brain-Computer-Interface development (Onose et al., 2012; Ivorra et al., 2018).
Especially reading task intention from gaze allows to use low-level cognitive inputs to realize systems with high autonomy, making it easier to interact with the system (González-Díaz et al., 2024). In contrast to these possibilities, there are some challenges that need to be considered when implementing gaze as a control input:
@@
  • Unintended selection: When gaze is used as a pointing device, e.g. to select an object, rules must be applied to ensure, that not all fixations are selection triggers. This issue is called the Midas Touch problem and describes an unintentional selection of an object (Dünser et al., 2015; Jones et al., 2018; Stalljann et al., 2020). Reasons include unintentional and reflexive eye movements, overlapping objects, distraction, stress, and limited accuracy of the eye-tracking device (Holmqvist and Andersson, 2017).
@@
  • Context misinterpretation: It is caused by overlapping objects or task ambiguity. When objects are close together and the eye-tracker accuracy is not sufficient, the gaze can be tracked on the wrong object leading to faulty interactions between robot and object (Jones et al., 2018). In the interview conducted by Herr et al. (Herr et al., 2016) participants stated: “Well, now I look at you [the interviewer], but I actually focus on the bottle behind you. That would be difficult for the robot to recognize”. This behavior suggests that attention to a particular location is sometimes unrelated to intention or cognitive process.
  • Physiological factors: Physiologically, there are some factors that affect the accuracy of the gaze. Neurological conditions, fatigue, drooping eyelids and strabismus affect the functionality of eye-tracking devices (Holmqvist and Andersson, 2017). The robotic workspace and gaze movements differ in dimensions. Three-dimensional robot workspace has to be covered by two-dimensional eye movements. Eye tracking glasses have decreasing accuracy in the periphery of the field of view. When head movement is limited, as in the case of the stakeholder group, these peripheral regions are more likely to be used to fixate objects, leading to increased instances of contextual misinterpretation.
@@
When such challenges occur, it can lead to frustration and high cognitive load for the user to avert such situations (Dünser et al., 2015; Jones et al., 2018).

### v2

@@ Abstract.
Eye-tracking-driven controls for assistive robotic arms provide people with severe physical disabilities with intuitive interaction opportunities.In this context, shared control can improve acceptance of the robot by partially automating task execution. Based on recent literature, we present a Wizard of Oz design for shared control driven by eye-tracking. This approach allows for the rapid exploration of user expectations to inform future design iterations. Two studies were conducted to assess user experience, identify design challenges, and determine ways to improve usability and accessibility. The first study involved people with severe motor disabilities (PSMD) and consisted of an online survey. The second study aimed to gain technical insights through direct robot interaction to provide a comprehensive overview of the findings.
@@
copyright: rightsretainedjournal: THRIjournalyear: 2026journalvolume: 1journalnumber: 1publicationmonth: 1doi: 10.1145/3796524journal: JACMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Human-centered computing Accessibility systems and toolsccs: Human-centered computing Pointing devicesccs: Human-centered computing Empirical studies in accessibilityccs: Software and its engineering Development frameworks and environments
@@ 1. Introduction
People with severe motor disabilities (PSMD) often rely on hands-free control modalities, such as Brain-Computer Interfaces, sip-and-puff systems, and eye-tracking interfaces. These technologies have been successfully used by individuals with conditions such as locked-in syndrome, cerebral palsy, late-stage amyotrophic lateral sclerosis, or multiple sclerosis. As control inputs for assistive robotic arms, they can support independence, self-determination, and privacy in daily life, reducing reliance on family members and caregivers (Fischer-Janzen et al., 2024; Awuah et al., 2024; Baumeister et al., 2021).
A shared control allows users to give the robot high-level commands, such as selecting an object or goal, by automating task sequences. In eye-tracking-driven approaches, advances in AI allow us to move beyond manual mappings toward intelligent interfaces that interpret gaze patterns to extract task intent (Aronson et al., 2018; Belardinelli, 2024; González-Díaz et al., 2024). Environmental context, such as the presence of objects, can further improve autonomous decision-making (Zhang et al., 2025; Yang et al., 2021). These advances have led to the development of various robot control methods for Human-Robot Interaction (HRI) that differ in terms of control strategies (Mohebbi, 2020; Fischer-Janzen et al., 2024). With a higher level of robot autonomy, challenges such as controlling three-dimensional robot movement with two-dimensional eye movements can be resolved. Bypassing the need for mode switches, as required for direct control, to cover the 3-dimensional robot space, or independent joint control, can reduce workload and decrease the concentration required for the task (Fischer-Janzen et al., 2024; Kronhardt et al., 2023; Herlant et al., 2016). Shared control approaches have been shown to positively impact interaction perception and reduce task completion time (Bhattacharjee et al., 2020; Kronhardt et al., 2023). A challenge associated with a high level of automation is maintaining the sense of agency. Community feedback is essential to improve design and evaluate usability and accessibility (Stalljann et al., 2020; Baumeister et al., 2021; Nanavati et al., 2024a), as well as to uncover influences and design controls with an appropriate level of autonomy. It provides firsthand insights into user expectations, thus enhancing acceptance rates (Nanavati et al., 2024a; Canal et al., 2021).
The novelty of this work is characterized by a detailed look at usability and functionality regarding robot acceptance parameters in respect to articulated robots used to assist in daily life. Current task intent recognition methods are based on object labeling done by object detection or segmentation, which we evaluated in this context on feasibility. This work incorporates community feedback on an initial shared control design to identify challenges and guide improvements. A Wizard of Oz design of the eye-tracking-driven control was developed and used in two studies. The first study was an online survey distributed to PSMD, their families, friends, and medical professionals. The goal of the study was to assess everyday usability and identify task preferences. The stakeholder group was defined as PSMD who experience significant motor impairments in their arms and torso. Previous research has shown differences in the perception of virtual versus real robot interactions (Petrich et al., 2022). A second study was conducted, in which participants without motor impairments and with varying levels of robotics experience interacted with the physical system. This study focused on deriving technical characteristics that can inform the constraints of a future control algorithm, such as feasible dwell times and accuracy requirements. The main contributions of this work are as follows:
@@
    Presentation of an initial eye-tracking driven shared control design leveraging object detection algorithms for daily tasks.
@@
    A functionality and usability evaluation, incorporating feedback from potential users, their families, medical professionals, and non-disabled participants with and without prior robot experience.
@@
    Refinements and task preferences for the initial shared control design, derived from insights obtained in both studies.
@@ 2. Background
2.1. Gaze as control input in shared controls
@@
”Shared control in assistive robotics blends human autonomy with computer assistance, thus simplifying complex tasks for individuals with physical impairments” (Goldau et al., 2024). Such a system can achieve lower performance times and error rates when the user provides strategic intent and the autonomy supplies continuous, low-level stabilization, obstacle avoidance, and motion refinement (Farhadi et al., 2025). Advances in sensing, inference, modeling, and learning methods have led to novel developments in this research area (Selvaggio et al., 2021).
Gaze can be used as an implicit or explicit control input for shared controls (Loke et al., 2025). As an implicit control input, natural gaze patterns are used to recognize or predict user intent. For example, gaze behaviors such as fixating on task-relevant objects before naming or manipulating them can be used and predicted before the desired task even begins (Admoni and Scassellati, 2017; Aronson and Admoni, 2020; Belardinelli, 2024; González-Díaz et al., 2024; Selvaggio et al., 2021). Huang and Mutlu demonstrated an example of such controls, in which ingredients were predicted by gaze patterns (Huang and Mutlu, 2016). Admoni et al. created a shared control system based on a POMDP (Partially Observable Markov Decision Process), which predicts single tasks from gaze patterns in an unstructured environment (Admoni and Srinivasa, 2016). Zhang et al. presented another approach based on the usage of a large language model (LLM) and vision foundation model (VFM), that recognizes user intent by interpreting labels created by the VFM and using the LLM to find common object combinations (Zhang et al., 2025). Wang used machine learning to recognize actions based on 3D gaze-related features to control a robot (Wang, 2021). Lastly, Farhadi et al. presented a review of BCI-related controls and shared control approaches (Farhadi et al., 2025). Challenges in predicting patterns of natural gaze can arise from multiple cognitive processes occurring in our brain at any time. As Belardinelli states, this issue worsens when performing activities more complex than grasping or selecting an object (Belardinelli, 2024). If incorrect or imprecise predictions are made, collisions or large penalties in completion time and user preferences can occur (Jain and Argall, 2019). As Cao et al. point out, timely and accurate estimation of intentions and, consequently, prediction remain major challenges (Cao et al., 2025).
@@
Gaze, as an explicit input, can vary in its function in control. The first eye-tracking controllers for robotic arms were developed as early as 2001 (Kim et al., 2001). Since then, several systems have been developed, but only a few have become commercially available (Fischer-Janzen et al., 2024; Suryadarma et al., 2024; Brose et al., 2010). One example of explicit control input is directional control, in which the robot moves in the direction of gaze (Alsharif, 2017; Sunny et al., 2021). Graphical User Interfaces (GUIs) are available in which the robot is moved with gaze-selected buttons, taking advantage of more accurate stationary eye-tracking systems (Dragomir et al., 2021; Stalljann et al., 2020). Such controls are easy to learn, provide a good sense of agency, and reduce cognitive load (Dragomir et al., 2021; Dziemian et al., 2016). However, controlling each of the robot movements separately is more time consuming. A popular approach, as outlined in Belardinelli’s work, is to use relevant objects in the scene and define them as Areas-Of-Interest (AOI). Fixations within these areas can be evaluated and, together with a certain dwell time, used as a selection method (Catalán et al., 2017; Zeng et al., 2017; Tostado et al., 2016). We refer to the work of Fischer-Janzen et al. and Suryadarma et al. for examples in realization, due to the number of related works (Fischer-Janzen et al., 2024; Suryadarma et al., 2024). A major limitation of these works is that they do not allow for the subsequent selection of multiple tasks, which is a major requirement for daily use, as demonstrated by Hagengruber et al. (Hagengruber et al., 2025). In this work, gaze is used as explicit input by focusing on the desired object.
Challenges can occur in both implicit and explicit control inputs that need to be considered:
@@
  • Unintended selection: When the gaze is used as a pointing device, e.g., to select an object, rules must be applied to ensure that not all fixations do not trigger selections. These unintentional selections are described by the Midas Touch problem (Dünser et al., 2015; Jones et al., 2018; Stalljann et al., 2020). Reasons for this problem include unintentional and reflexive eye movements, overlapping objects, distraction, stress, and the limited accuracy of the eye-tracking device (Holmqvist and Andersson, 2017). When objects are close together and the eye tracker’s accuracy is insufficient, the gaze may be tracked to the wrong object (Jones et al., 2018).
@@
  • Context misinterpretation: The aforementioned parallel cognitive processes can result in task ambiguity. In an interview conducted by Herr et al. (Herr et al., 2016), participants stated: “Well, now I look at you [the interviewer], but I actually focus on the bottle behind you. That would be difficult for the robot to recognize”. This behavior suggests that attention to a particular location is sometimes unrelated to intention or cognitive process.
  • Physiological factors: There are some physiological factors that affect the accuracy of the gaze. Neurological conditions, fatigue, drooping eyelids, and strabismus can affect the functionality of eye-tracking devices (Holmqvist and Andersson, 2017).
@@
Such challenges can lead to frustration and a high cognitive load for users trying to avoid them (Dünser et al., 2015; Jones et al., 2018).
2.2. Characteristics in gaze-based shared controls
Our system design was inspired by the impact factors on user acceptance of shared control found in the current literature. Many characteristics and technology acceptance models exist (Beer and Akanksha Prakash, 2011). Due to the multifaceted nature of user acceptance, we present only the factors that influenced our design. For more information on characteristics influencing user acceptance, see the works of Smarr et al., Rahman et al., Riek et al., and Beer et al. (Smarr et al., 2012; Rahman et al., 2025; Riek and Howard, 2014; Beer and Akanksha Prakash, 2011). We introduce three categories to facilitate an overview of the results: Technical challenges, level of autonomy, and user preferences.
Technical challenges include ensuring the system’s reliability and functionality. The system must be robust in changing or unorganized environments, especially when the robot is mounted on an electric wheelchair (Selvaggio et al., 2021; Stanger et al., 1994). The system’s capabilities related to task performance also define its functionality in everyday life. We discussed such challenges in Section 2.1. The subtasks necessary to correctly complete a task must be understood and set in the proper context relative to each other.
The level of autonomy is a subject of ongoing debate and can range from tele-operated systems to fully autonomous robots (Beer et al., 2014). Recent studies have shown that more autonomy does not necessarily align with user preferences (Nanavati et al., 2024b; Bhattacharjee et al., 2020). If a robot with a high level of autonomy produces more errors in task execution, users tend to reject the system (Bhattacharjee et al., 2020). This can lead to a loss of trust or a reduced sense of agency due to the system’s unpredictable behavior (Nanavati et al., 2024b; Baptista et al., 2025). Losey et al. summarized this as a duality: More automation reduces the user’s burden when performing tasks, but it also limits the mutual support when it ignores the user’s intent. The autonomous behavior reduces the operator’s workload when performing repetitive tasks (Selvaggio et al., 2021). As Loke et al. state, high mental workload contributes to the abandonment of assistive devices (Loke et al., 2025). Due to the potential shift in cognitive resources toward the task and away from precise robot movements, control can benefit from being easy to understand (Mazzola et al., 2025; Selvaggio et al., 2021). However, this may vary depending on the complexity of the task (Stanger et al., 1994; Shafti et al., 2019; Goldau et al., 2024), leading back to the initial question of what level of autonomy is most efficient and reliable in assisting the user.
Users’ preferences vary based on personal metrics such as the user’s age, level of disability, living arrangements, and familiarity with technology (Stanger et al., 1994; Smarr et al., 2012; Nanavati et al., 2024b). Incorporating user preferences influences perception of and satisfaction with the system (Cakmak et al., 2011; Canal et al., 2021). Parameters such as speed can influence the user perception (Bhattacharjee et al., 2020). Another example is the participant’s preference for handing over an object with a particular orientation (Cakmak et al., 2011). The ability to adapt the system seamlessly to users’ preferences regarding the level of assistance has given rise to shared autonomy approaches. Those are based on the autonomy’s understanding of the user’s intentions and of the surrounding environment (Selvaggio et al., 2021). In this context, research focuses on arbitration strategies that adapt the level of autonomy to the user’s preferences, as demonstrated in the works of Dragan et al., Losey et al., and Jain et al. (Dragan and Srinivasa, 2013; Jain and Argall, 2019; Losey et al., 2018). The arbitration scheme, together with the interaction paradigm, describes the collaboration between human and robot (Farhadi et al., 2025).
When designing well-balanced shared control, a trade-off must be found for all technical and ethical factors. Section 3.1 presents a solution that considers and balances the different influences.
2.3. Tasks in assistive robot manipulation
Edemekong et al. describe Activities of Daily Living [(ADLs)] as basic routine tasks that most healthy individuals can perform without assistance (Edemekong et al., 2025). ADLs are often used to evaluate a person’s level of disability, as in the WHODAS 2.0 (World Health Organization, 2010). Nanavati et al. used ADLs in their review to compare the responses of individuals requiring ADL assistance with the work of researchers in the field of assistive robotics. A major misalignment was found. While most researchers focus on mobility and eating, people in need of assistance state that they need help with going out, managing money, doing housework, preparing meals, bathing, and dressing (Nanavati et al., 2024b). Petrich et al. took a different approach, examining the behavior of people in everyday life and uncovering results quite different from those previously mentioned. In their study, tasks such as reaching to pick up and place items, opening and closing doors, and using switches and buttons were most important, according to video analysis. These results differed from the participants’ survey answers (Petrich et al., 2022). These results highlight that different approaches can reveal various aspects of task prioritization and preferences in HRI. Smarr et al. explored robot acceptance among the elderly and found that the elderly accept robots more readily when they assist with learning new skills and hobbies. However, human support was preferred for communication, which emphasizes the importance of social contact (Smarr et al., 2012). The examples presented above show that assistive robotic arms cannot directly assist with activities such as bathing or managing finances. It is important to distinguish between robotic arm-related tasks and others.
As stated by Hagengruber et al., the purpose of assistive robotics is to carry out various tasks in succession in order to depict a realistic daily routine (Hagengruber et al., 2025). They presented an sEMG-based control system that could perform a drinking task, including retrieving a cup from a drawer and pouring a drink. Chi et al. demonstrated that dividing complex tasks into subtasks facilitates accomplishment. For example, they divided an eating task into subtasks such as getting a spoon, scooping food, and bringing it near the user’s mouth (Chi et al., 2025). Using gaze as a control input, one can obtain gaze patterns specific to each subtask that can be used to train AI models (Hayhoe and Ballard, 2014).
Robots can perform daily tasks, such as household chores and personal hygiene, as well as medical procedures (Herr et al., 2016; Rahman et al., 2025; Bien et al., 2003; Rahman et al., 2025; Wang et al., 2020). The preferred tasks vary depending on different influences, such as the user’s attitude towards a task. Delegating household tasks to robots is one example of this (Herr et al., 2016). Similarly, Petrich et al. suggested automating tasks that are not prioritized by the community to reduce cognitive load (Petrich et al., 2022).
Lastly, one must determine whether a robotic arm is the right tool for the task at hand. For example, Petrich et al. stated that robots can help with using electronics, such as smartphones; however, it is more convenient to use a hands-free phone (Petrich et al., 2022). Similar results were found for toileting, where people preferred the aid of caregivers and smart home applications. Hagengruber et al.’s work, in particular, demonstrates the advantage of using multiple assistive devices together. For example, it simplified opening a drawer by holding the handle with the robot and moving the wheelchair backward (Hagengruber et al., 2025).
In addition to the work of Nanavati et al. (Nanavati et al., 2024b), our review revealed that the four most frequently mentioned tasks, suitable for assistive robots, were food related tasks (serving a meal, eating) (Bien et al., 2003; Stanger et al., 1994; Jardón et al., 2012; Chang and Park, 2003; Canal et al., 2021; Bhattacharjee et al., 2020; Nanavati et al., 2024a), drinking related tasks (Bien et al., 2003; Stanger et al., 1994; Jardón et al., 2012; Chang and Park, 2003), picking up objects (Bien et al., 2003; Stanger et al., 1994; Jardón et al., 2012; Chang and Park, 2003; Nanavati et al., 2024b), and personal hygiene (Stanger et al., 1994; Jardón et al., 2012; Nanavati et al., 2024b). Personal hygiene included various tasks such as shaving (Bien et al., 2003; Chang and Park, 2003), washing the face and hands, brushing teeth, combing hair, shaving, applying makeup (Jardón et al., 2012), and dressing (Canal et al., 2021; Nanavati et al., 2024b). The importance of being able to manipulate a wide variety of objects benefits the ability to get around, e.g., with an electric wheelchair, by being able to move blocking objects out of the way (Nanavati et al., 2024b).
Other miscellaneous tasks were mentioned with taking medicine (Nanavati et al., 2024b), working (Nanavati et al., 2024b), leisure activities (Stanger et al., 1994; Nanavati et al., 2024a), turning on and off switches (Bien et al., 2003; Chang and Park, 2003), opening and closing doors (Bien et al., 2003; Brose et al., 2010; Chang and Park, 2003), scratching oneself (Bien et al., 2003; Chang and Park, 2003), and making tea. Sources from 2003 show that changing CDs and removing paper from fax machines are examples of how task preferences change over time (Bien et al., 2003; Chang and Park, 2003). As the accessibility of smart home applications continues to evolve, it is expected that more tasks will become less important for robot assistance, such as flipping light switches or opening doors.
Based on these results, tasks were selected for the Wizard of Oz design. In Section 4.1, the literature findings were used to group the responses from the online survey. In Section 5.1, they were used to compare the results in terms of changes in task prioritization over time.